# Stochastic asymptotical regularization for nonlinear ill-posed problems

### v2

@@ Abstract
Recently, the stochastic asymptotical regularization (SAR) has been developed in (Inverse Problems, 39: 015007, 2023) for the uncertainty quantification of the stable approximate solution of linear ill-posed inverse problems. In this paper, we extend the regularization theory of SAR for nonlinear inverse problems. By combining techniques from classical regularization theory and stochastic analysis, we prove the regularizing properties of SAR with regard to mean-square convergence. The convergence rate results under the canonical sourcewise condition are also studied. Several numerical examples are used to show the accuracy and advantages of SAR: compared with the conventional deterministic regularization approaches for deterministic inverse problems, SAR can quantify the uncertainty in error estimates for ill-posed problems, improved accuracy by selecting the optimal path, escape local minimums for nonlinear problems, and identify of multi-solutions by clustering samples of obtained approximate solutions.
@@ 1 Introduction
The continuous analog to (1.5) as $\Delta t$ tends to zero is known as asymptotic regularization or Showalter’s method (see, e.g., [6]). It is written as a first-order evolution equation of the form
@@
This continuous framework has also received much attention in the development of modern regularization theory of inverse problems. For instance, the authors in [11, 13, 12, 14] extended (1.4) to second-order and fractional-order gradient flows, which exhibit acceleration affect. The authors in [9] studied an asymptotical regularization with convex constraints for nonlinear inverse problems with special features of solutions, such as sparsity and discontinuity. Furthermore, by connecting to some classical methods (i.e. the Kalman-Bucy filter and 3DVAR) from data assimilation, the authors in [10] studied stationary and non-stationary asymptotical regularization methods for linear inverse problems (1.1) in the presence of white noise.
@@
where the auxiliary function $f(t)\in L^{\infty}({\mathbb{R}}_{+})$ and $B_{t}$ is an ${\mathcal{X}}$-valued $Q$-Wiener process, is also gain much attention in the field of optimization and marching learning, see [19, 20] and references therein.
Recently, motivated from the uncertainty quantification for infinite dimensional inverse problems, the Stochastic Asymptotical Regularization (SAR) method, i.e. the linear counterpart of (1.6), had been introduced in [21, 22] for linear inverse problems. It has been show that compared with conventional deterministic regularization approaches for deterministic inverse problems such as (1.1)-(1.2), SAR exhibits four new capabilities: (i) it can quantify the uncertainty in error estimates for inverse problems (e.g. the a posteriori (pointwise) error estimation, etc.), see [21, Section 2.2] for a rigorous mathematical description. (ii) It can reveal and explicate hidden information for many real-world problems that are usually obscured by incomplete mathematical modeling and the presence of large-scale noise, see [21, Section 5.2] for an example in biosensor tomography. (iii) It can improved accuracy by selecting the optimal path [22, Section 4.1], and (iv) it can identify multi-solutions by clustering samples of obtained approximate solutions [22, Section 4.3]. Moreover, in comparison with the Bayesian methods, SAR requires neither noise structure nor a prior probability distribution of the exact solution. Building on this idea, in this paper we aim to extend SAR for non-linear inverse problems (1.1)-(1.4). In this paper, we will prove that the stochastic differential equation (1.6), associated with an appropriate a posteriori stopping rule, yields a regularization method, also called the Stochastic Asymptotical Regularization (SAR).
@@ 2.1 Assumptions
Assumption 2.1(i) imposes boundedness and continuity on the derivative $F^{\prime}(x)$. The inequality (2.2) is often known as the tangential cone condition, and it controls the degree of nonlinearity of the operator $F$. Roughly speaking, it requires that the forward map $F$ be not too far from a linear map. Moreover, since the solution to problem (1.6) may be nonunique, in this paper the reference solution $x^{\dagger}$ is taken to be the minimum norm solution (with respect to the initial guess $\bar{x}$), which is known to be unique under Assumption 2.1(ii) (see, e.g., [23]).
@@
Under Assumption 2.1, the following statements hold true:
@@
Also assume that, for all $x\in B_{r}(\bar{x})$, there exist a linear bounded operator $R_{x}:\mathcal{Y}\rightarrow\mathcal{Y}$ and a constant $c_{R}\geq 0$ such that
@@ 2.2 Necessary concepts of stochastic calculus
3 Convergence
Throughout this paper, the notation $c$, with a subscript or/and superscript, denotes a generic constant, which may differ at each occurrence but is always independent of the noise level $\delta$ and the time variable $t$. An example of these constants can be found in Table LABEL:NotationTable in Appendix D. The notation $(\mathbb{E}[X^{a}])^{b}$ will be written as $\mathbb{E}[X^{a}]^{b}$, with $a,b>0$ for simplification.
@@ 3 Convergence
Let $x^{*}$ and $x^{\delta}(t)$ be the solutions to problems (1.1) and (1.6), respectively. If Assumption 2.1 is satisfied, the following holds:
@@
Consequently, by taking the derivative of (3.11) with respect to $t$, we get
@@
Let $x^{\delta}(t)$ be the solution to (1.6) and $x^{*}$ a solution to (1.1). Assume that ${\|{F\left({{\bar{x}}}\right)-{y^{\delta}}}\|}>\tau\delta$. Then, under Assumption 2.1, there exists a unique $t^{*}$, defined by (3.13).
@@
We can now state the convergence of SAR for the exact data $y$, provided Assumption 2.1 holds.
@@
Let $x(t)$ be the solution to (1.6), with exact data $y$ and $x^{*}$ a solution to (1.1). If Assumption 2.1 is fulfilled, the following holds:
@@
and show that $\mathbb{E}[\|e(t)\|^{2}]=0$ as $t\to\infty$:
@@
Let Assumption 2.1 be satisfied and ${\|{F\left({{\bar{x}}}\right)-{y^{\delta}}}\|}>\tau\delta$. Then, if the terminating time $t^{*}=t^{*}(\delta,y^{\delta})$ is determined by (3.13), there exists a solution $x^{*}$ to problem (1.1) such that
@@
Using Theorem 3.7, we can fix a large $\overline{T}$ so that the term $\mathbb{E}[{{{\|{{x}(t^{\delta}_{\overline{T}})-{x^{*}}}\|}^{2}}}]$ is sufficiently small. As $t^{\delta}_{\overline{T}}$ is fixed, we can safely say that ${x^{\delta_{{T}}}}(t^{\delta}_{\overline{T}})$ depends continuously on $y^{\delta}$ because of the representation of $x^{\delta}$ in (3.6) and Assumption 2.1(i). We can then conclude that $\mathbb{E}[{{{\|{{x^{\delta_{{T}}}}(t^{\delta}_{\overline{T}})-{x}(t^{\delta}_{\overline{T}})}\|}^{2}}}]$ must go to zero as $T\to 0$. This completes the proof of the first assertion. The case $\mathcal{N}(F^{\prime}(x^{\dagger}))\subset\mathcal{N}(F^{\prime}(x(t)))$ follows in a manner similar to that of Theorem 3.7. ∎
@@ 4.1 Estimation of the bias error
Let Assumptions 2.1 and 2.2 be satisfied. Choose $\epsilon_{0}\in(\max\{\frac{2-\sqrt{2}}{\eta}-(2+\sqrt{2}),0\},2(\frac{1}{\eta}-1))$. Then, for any $\sigma\geq 0$ , the following holds:
@@ 5.1 Model problem 1: a parameter-identification problem in PDEs
To use SAR in practice, we need to discretize the stochastic flow (1.6) with respect to the artificial time variable $t$. Numerous algorithms have been proposed (see, e.g., [33, 34]) for accurate numerical approximation of stochastic differential equations. However, for simplicity, here we consider only the Euler method, i.e.
@@
where $\theta$ is employed to describe the level of randomization, which ranges between 00 and $\frac{\sqrt{\epsilon_{0}\eta}}{\delta_{0}}$. The higher the value of $\theta$, the more randomization is assumed. For $\theta=0$, there will be no randomization, and this happens to be the conventional asymptotic regularization method as introduced in [6].
@@ 5.1.1 Settings for simulations
Unless otherwise stated, the examples are discretized with a dimensionality $n=256$ for 1D and $n^{2}=128^{2}=16384$ for 2D. The noisy data $u^{\delta}$ is generated from the exact data $u$ as follows:
@@
where $\delta$ is the noise level, and the random variables $\xi_{j}$ follow the standard Gaussian distribution. To fairly evaluate our method, we randomly select the initial guess $x_{0}$ in each trial of the numerical experiments. Furthermore, to assess the accuracy of the approximate solutions, we define the root mean-square error (RMSE) and the root mean-square residual (RMSR):
$\textmd{RMSE}:=\mathbb{E}[\|c^{\dagger}-c^{\delta}_{k}\|^{2}]^{\frac{1}{2}},\quad\textmd{RMSR}:=\mathbb{E}[\|F(c^{\delta}_{k})-u^{\delta}\|^{2}]^{\frac{1}{2}}.$
It is important to note that the level of randomization, $\theta$, and the sample size influence the behavior of SAR. A detailed numerical illustration of this effect can be found in Appendix E and Tables 1 and 2 below. With that in mind, let us investigate some advantages of SAR.
@@ 5.1.2 Advantage I: Escaping from local minima by selecting the optimal path
Additionally, Table 1 shows that with an appropriate level of randomization and a suitable sample size, the SAR method significantly improves the inversion results compared to the case of $\theta=0$ (Landeweber method). All relevant quantities, such as MinNorm, MeanNorm, $\mathbb{E}(n_{*})$, Min$(n_{*})$ and Med$(n_{*})$, are noticeably reduced.
@@
Figure 2: Estimation of $P(\|x_{n}-x^{\dagger}\|\geq\delta_{i})$ using $N=500$ independent runs for 1D case with $\theta=0.5$ and different noise levels $\delta_{1}=1\%,\delta_{2}=3\%,\delta_{3}=5\%$.
@@ 5.2 Model problem 2: the autoconvolution equation
$D^{\varepsilon}_{KL}(\mathbf{u}\|\mathbf{v}):=\sum_{i}|\mathbf{u}_{i}|\ln\left(\frac{|\mathbf{u}_{i}|+\varepsilon}{|\mathbf{v}_{i}|+\varepsilon}\right).$
@@
At the end of this section, we remark that it is possible to find another group of two exact functions corresponding to the obtained approximate solutions displayed in Figure 8. However, any two groups of exact functions exhibit similar features. Moreover, in most real-world problems, we do not know how many exact sources exist and what form they have. In such cases, it is sufficient to find some approximate solutions with significantly distinguished features, i.e., people mainly interest in the interpretation of a physical phenomenon by studying the properties of reconstructed solutions.
@@
6 Conclusions
We develop a statistical approach – SAR for deterministic nonlinear inverse problems. In this paper, a complete regularization theory of SAR has been presented and analyzed. In particular, it has been shown that SAR yields an optimal regularization method with regard to mean-square convergence. Compared with the conventional deterministic regularization method, SAR has three advantages: first, it can provide the uncertainty quantification of the estimated quantity, which is described by a deterministic forward model; second, SAR can escape from the local minima numerically, which is the main difficulty for most nonlinear inverse problems; third, SAR can identify multi-solutions by clustering solution samples. SAR will therefore be a useful tool for studying practical nonlinear inverse problems with deterministic ill-posed forward models.
@@ 6 Conclusions
It follows from Lemma 4.3 and Assumption 2.2 that
@@
which further implies, together with the choice of $\epsilon_{0}$ and Assumption 2.2, that
@@
$\begin{split}G_{1}(t,z_{1},z_{2})=&\left(1+\sqrt{\frac{4^{\gamma}-1}{\gamma}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma}}+{c_{1}}\sqrt{t+1}\cdot\frac{c^{*}E}{(1+t)^{\gamma+\frac{1}{2}}}\\ &+{c_{2}}\int_{0}^{t}{\frac{{\frac{c^{*}E}{(1+s)^{\gamma}}\cdot\frac{c^{*}E}{(1+s)^{\gamma+\frac{1}{2}}}}}{{\sqrt{1+t-s}}}}\textmd{ds}\\ \leq&\left(1+\sqrt{\frac{4^{\gamma}-1}{\gamma}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma}}+{c_{1}}c^{*}\frac{E}{(1+t)^{\gamma}}\\ &+c_{2}(c^{*})^{2}E\frac{4^{\gamma}-1}{\gamma}\frac{E}{(1+t)^{\gamma}}\\ =&\left(c^{b}_{\gamma}+\sqrt{\frac{4^{\gamma}-1}{\gamma}}c_{3}c^{b}_{\gamma}+{c_{1}}c^{*}+c_{2}(c^{*})^{2}\frac{4^{\gamma}-1}{\gamma}E\right)\frac{E}{(1+t)^{\gamma}},\end{split}$
@@
$\begin{split}G_{2}(t,z_{1},z_{2})=&\left(1+\sqrt{\frac{2^{2\gamma+2}-2}{2\gamma+1}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}+{c_{1}}\frac{c^{*}E}{(1+t)^{\gamma+\frac{1}{2}}}\\ &+{c_{2}}\int_{0}^{t}{\frac{{\frac{c^{*}E}{(1+t)^{\gamma}}\cdot\frac{c^{*}E}{(1+t)^{\gamma+\frac{1}{2}}}}}{{{1+t-s}}}}\textmd{ds}\\ \leq&\left(1+\sqrt{\frac{2^{2\gamma+2}-2}{2\gamma+1}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}+{c_{1}}c^{*}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}\\ &+c_{2}(c^{*})^{2}E\frac{2^{2\gamma+\frac{5}{2}}-4}{4\gamma+1}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}\\ =&\left(c^{b}_{\gamma}+\sqrt{\frac{2^{2\gamma+\frac{5}{2}}-4}{4\gamma+1}}c_{3}c^{b}_{\gamma}+{c_{1}}c^{*}+c_{2}(c^{*})^{2}E\frac{2^{2\gamma+\frac{5}{2}}-4}{4\gamma+1}\right)\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}.\end{split}$

### v3

@@ Abstract
Recently, the stochastic asymptotical regularization (SAR) has been developed in (Inverse Problems, 39: 015007, 2023) for the uncertainty quantification of the stable approximate solution of linear ill-posed inverse problems. In this paper, we extend the regularization theory of SAR for nonlinear inverse problems. By combining techniques from classical regularization theory and stochastic analysis, we prove the regularizing properties of SAR with regard to mean-square convergence. The convergence rate results under the canonical sourcewise condition are also studied. Several numerical examples are used to show the accuracy and advantages of SAR: compared with the conventional deterministic regularization approaches for deterministic inverse problems, SAR can quantify the uncertainty in error estimates for ill-posed problems, improve accuracy by selecting the optimal path, escape local minima for nonlinear problems, and identify multiple solutions by clustering samples of obtained approximate solutions.
@@ 1 Introduction
The continuous analog to (1.3) as $\Delta t$ tends to zero is known as asymptotical regularization or Showalter’s method (see, e.g., [6]). It is written as a first-order evolution equation of the form
@@
This continuous framework has also received much attention in the development of modern regularization theory of inverse problems. For instance, the authors in [11, 13, 12, 14] extended (1.4) to second-order and fractional-order gradient flows, which exhibit acceleration effect. The authors in [9] studied an asymptotical regularization with convex constraints for nonlinear inverse problems with special features of solutions, such as sparsity and discontinuity. Furthermore, by connecting to some classical methods (i.e. the Kalman-Bucy filter and 3DVAR) from data assimilation, the authors in [10] studied stationary and non-stationary asymptotical regularization methods for linear inverse problems (1.1) in the presence of white noise.
@@
where the auxiliary function $f(t)\in L^{\infty}({\mathbb{R}}_{+})$ and $B_{t}$ is an ${\mathcal{X}}$-valued $Q$-Wiener process, has also gained much attention in the field of optimization and machine learning, see [19, 20] and references therein.
Recently, motivated by the uncertainty quantification for infinite dimensional inverse problems, the Stochastic Asymptotical Regularization (SAR) method, i.e. the linear counterpart of (1.6), had been introduced in [21, 22] for linear inverse problems. It has been shown that compared with conventional deterministic regularization approaches for deterministic inverse problems such as (1.1)-(1.2), SAR exhibits four new capabilities: (i) it can quantify the uncertainty in error estimates for inverse problems (e.g. the a posteriori (pointwise) error estimation, etc.), see [21, Section 2.2] for a rigorous mathematical description. (ii) It can reveal and explicate hidden information for many real-world problems that are usually obscured by incomplete mathematical modeling and the presence of large-scale noise, see [21, Section 5.2] for an example in biosensor tomography. (iii) It can improve accuracy by selecting the optimal path [22, Section 4.1], and (iv) it can identify multi-solutions by clustering samples of obtained approximate solutions [22, Section 4.3]. Moreover, in comparison with the Bayesian methods, SAR requires neither noise structure nor a priori probability distribution of the exact solution. Building on this idea, in this paper we aim to extend SAR for non-linear inverse problems (1.1)-(1.4). In this paper, we will prove that the stochastic differential equation (1.6), associated with an appropriate a posteriori stopping rule, yields a regularization method, which we refer to Stochastic Asymptotical Regularization (SAR).
@@ 2.1 Assumptions
Assumption 1(i) imposes boundedness and continuity on the derivative $F^{\prime}(x)$. The inequality (2.2) is often known as the tangential cone condition, and it controls the degree of nonlinearity of the operator $F$. Roughly speaking, it requires that the forward map $F$ be not too far from a linear map. Moreover, since the solution to problem (1.6) may be nonunique, in this paper the reference solution $x^{\dagger}$ is taken to be the minimum norm solution (with respect to the initial guess $\bar{x}$), which is known to be unique under Assumption 1(ii) (see, e.g., [23]).
@@
Under Assumption 1, the following statements hold true:
@@
Also assume that, for all $x\in B_{r}(\bar{x})$, there exists a linear bounded operator $R_{x}:\mathcal{Y}\rightarrow\mathcal{Y}$ and a constant $c_{R}\geq 0$ such that
@@ 2.2 Necessary concepts of stochastic calculus
3 Convergence analysis
Throughout this paper, the notation $c$, with a subscript and/or superscript, denotes a generic constant, which may differ at each occurrence but is always independent of the noise level $\delta$ and the time variable $t$. An example of these constants can be found in Table LABEL:NotationTable in Appendix D. The notation $(\mathbb{E}[X^{a}])^{b}$ will be written as $\mathbb{E}[X^{a}]^{b}$, with $a,b>0$ for simplification.
@@ 3 Convergence analysis
Let $x^{*}$ and $x^{\delta}(t)$ be the solutions to problems (1.1) and (1.6), respectively. If Assumption 1 is satisfied, then the following inequality holds:
@@
Differentiating (3.11) with respect to $t$, we get
@@
Let $x^{\delta}(t)$ be the solution to (1.6) and $x^{*}$ a solution to (1.1). Assume that ${\|{F\left({{\bar{x}}}\right)-{y^{\delta}}}\|}>\tau\delta$. Then, under Assumption 1, there exists a unique $t^{*}$, defined by (3.13).
@@
We can now state the convergence of SAR for the exact data $y$, provided Assumption 1 holds.
@@
Let $x(t)$ be the solution to (1.6), with exact data $y$ and $x^{*}$ a solution to (1.1). If Assumption 1 is fulfilled, the following holds:
@@
and will show that $\mathbb{E}[\|e(t)\|^{2}]=0$ as $t\to\infty$.
@@
Let Assumption 1 be satisfied and ${\|{F\left({{\bar{x}}}\right)-{y^{\delta}}}\|}>\tau\delta$. Then, if the terminating time $t^{*}=t^{*}(\delta,y^{\delta})$ is determined by (3.13), there exists a solution $x^{*}$ to problem (1.1) such that
@@
Using Theorem 3.7, we can fix a large $\overline{T}$ so that the term $\mathbb{E}[{{{\|{{x}(t^{\delta}_{\overline{T}})-{x^{*}}}\|}^{2}}}]$ is sufficiently small. As $t^{\delta}_{\overline{T}}$ is fixed, we can safely say that ${x^{\delta_{{T}}}}(t^{\delta}_{\overline{T}})$ depends continuously on $y^{\delta}$ because of the representation of $x^{\delta}$ in (3.6) and Assumption 1(i). We can then conclude that $\mathbb{E}[{{{\|{{x^{\delta_{{T}}}}(t^{\delta}_{\overline{T}})-{x}(t^{\delta}_{\overline{T}})}\|}^{2}}}]$ must go to zero as $T\to 0$. This completes the proof of the first assertion. The case $\mathcal{N}(F^{\prime}(x^{\dagger}))\subset\mathcal{N}(F^{\prime}(x(t)))$ follows in a manner similar to that of Theorem 3.7. ∎
@@ 4.1 Estimation of the bias error
Let Assumptions 1 and 2 be satisfied. Choose $\epsilon_{0}\in(\max\{\frac{2-\sqrt{2}}{\eta}-(2+\sqrt{2}),0\},2(\frac{1}{\eta}-1))$. Then, for any $\sigma\geq 0$ , the following holds:
@@ 5.1 Model problem 1: a parameter-identification problem in PDEs
To use SAR in practice, we need to discretize the stochastic flow (1.6) with respect to the artificial time variable $t$. Numerous algorithms have been proposed (see, e.g., [33, 34]) for accurate numerical approximation of stochastic differential equations. However, for simplicity, we focus only on the Euler method here, which coincides with the stochastic gradient descent algorithm mentioned, as seen in formula (1.5), i.e.
@@
where $\theta$ is employed to describe the level of randomization, which ranges between 00 and $\frac{\sqrt{\epsilon_{0}\eta}}{\delta_{0}}$. The higher the value of $\theta$, the more randomization is assumed. For $\theta=0$, there will be no randomization, and this happens to be the conventional asymptotical regularization method as introduced in [6].
@@ 5.1.1 Settings for simulations
Unless otherwise stated, these examples are discretized with a dimensionality $n=256$ for the 1D case and $n^{2}=128^{2}=16384$ for 2D. The noisy data $u^{\delta}$ is generated from the exact data $u$ as follows:
@@
where $\delta$ is the noise level, and the random variables $\xi_{j}$ follow the standard Gaussian distribution. To ensure a fair evaluation of our method, the initial guess $x_{0}$ is randomly selected in each trial of the numerical experiments.
It is crucial to note that the level of randomization, the parameter $\theta$, and the sample size significantly impact the behavior of SAR. A detailed numerical illustration of this effect can be found in Appendix E, along with Tables 1 and 2 below. With that in mind, let us investigate the advantages of SAR.
@@ 5.1.2 Advantage I: Escaping from local minima by selecting the optimal path
Additionally, Table 1 shows that with an appropriate level of randomization and a suitable sample size, the SAR method significantly improves the inversion results compared to the case of $\theta=0$ (Landweber method). All relevant quantities, such as MinNorm, MeanNorm, $\mathbb{E}(n_{*})$, Min$(n_{*})$ and Med$(n_{*})$, are noticeably reduced.
@@
Figure 2: Estimation of $P(\|x_{n}-x^{\dagger}\|\geq\delta_{i})$ using $N=500$ independent runs for 1D case with $\theta=0.5$ and different noise levels $\delta_{1}=1\%,\delta_{2}=3\%,\delta_{3}=5\%$. The numerical results are derived from 500 independent simulations, with the starting point uniformly sampled using the formula $0.01+4\cdot\text{rand}$.
@@ 5.2 Model problem 2: the autoconvolution equation
$D^{\varepsilon}_{KL}(\mathbf{u}\|\mathbf{v}):=\sum^{n}_{i=1}|[\mathbf{u}]_{i}|\ln\left(\frac{|[\mathbf{u}]_{i}|+\varepsilon}{|[\mathbf{v}]_{i}|+\varepsilon}\right)$
for two vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$. Here, $[\mathbf{u}]_{i}$ denotes the i-th component of vector $\mathbf{u}$.
@@
6 Conclusion and outlook
We develop a statistical approach – SAR for deterministic nonlinear inverse problems. In this paper, we present and analyze a comprehensive regularization theory of SAR. Specifically, we demonstrate that SAR provides an optimal regularization method with respect to mean-square convergence. Compared to conventional deterministic regularization methods, SAR offers three key advantages: first, it enables uncertainty quantification of the estimated quantity, which is modeled by a deterministic forward model; second, SAR can numerically avoid local minima, a common challenge in most nonlinear inverse problems; and third, SAR can identify multiple solutions by clustering solution samples. Thus, SAR serves as a valuable tool for studying practical nonlinear inverse problems with deterministic ill-posed forward models.
Based on the convergence analysis in Sections 3-4, proving the regularization properties of SAR for inverse problems with stochastic forward models is straightforward. However, a compelling question is how to design an efficient SAR with specific parameters that must be carefully selected according to the interplay of randomness among data noise, the forward model, and the method itself. This careful selection is essential to efficiently quantify uncertainties arising from the randomness of both the noise and the forward model.
@@ 6 Conclusion and outlook
It follows from Lemma 4.3 and Assumption 2 that
@@
which further implies, together with the choice of $\epsilon_{0}$ and Assumption 2, that
@@
$\begin{split}&G_{1}(t,z_{1},z_{2})\\ &\quad=\left(1+\sqrt{\frac{4^{\gamma}-1}{\gamma}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma}}+{c_{1}}\sqrt{t+1}\cdot\frac{c^{*}E}{(1+t)^{\gamma+\frac{1}{2}}}+{c_{2}}\int_{0}^{t}{\frac{{\frac{c^{*}E}{(1+s)^{\gamma}}\cdot\frac{c^{*}E}{(1+s)^{\gamma+\frac{1}{2}}}}}{{\sqrt{1+t-s}}}}\textmd{ds}\\ &\quad\leq\left(1+\sqrt{\frac{4^{\gamma}-1}{\gamma}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma}}+{c_{1}}c^{*}\frac{E}{(1+t)^{\gamma}}+c_{2}(c^{*})^{2}E\frac{4^{\gamma}-1}{\gamma}\frac{E}{(1+t)^{\gamma}}\\ &\quad=\left(c^{b}_{\gamma}+\sqrt{\frac{4^{\gamma}-1}{\gamma}}c_{3}c^{b}_{\gamma}+{c_{1}}c^{*}+c_{2}(c^{*})^{2}\frac{4^{\gamma}-1}{\gamma}E\right)\frac{E}{(1+t)^{\gamma}}\end{split}$
@@
$\begin{split}&G_{2}(t,z_{1},z_{2})\\ &\quad=\left(1+\sqrt{\frac{2^{2\gamma+2}-2}{2\gamma+1}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}+{c_{1}}\frac{c^{*}E}{(1+t)^{\gamma+\frac{1}{2}}}+{c_{2}}\int_{0}^{t}{\frac{{\frac{c^{*}E}{(1+t)^{\gamma}}\cdot\frac{c^{*}E}{(1+t)^{\gamma+\frac{1}{2}}}}}{{{1+t-s}}}}\textmd{ds}\\ &\quad\leq\left(1+\sqrt{\frac{2^{2\gamma+2}-2}{2\gamma+1}}c_{3}\right)c^{b}_{\gamma}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}+{c_{1}}c^{*}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}+c_{2}(c^{*})^{2}E\frac{2^{2\gamma+\frac{5}{2}}-4}{4\gamma+1}\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}\\ &\quad=\left(c^{b}_{\gamma}+\sqrt{\frac{2^{2\gamma+\frac{5}{2}}-4}{4\gamma+1}}c_{3}c^{b}_{\gamma}+{c_{1}}c^{*}+c_{2}(c^{*})^{2}E\frac{2^{2\gamma+\frac{5}{2}}-4}{4\gamma+1}\right)\frac{E}{(1+t)^{\gamma+\frac{1}{2}}}.\end{split}$